Find the speed of light in flint glass, water, and
Answer #1

speed of light in material = c/refractive index of material

a) speed of light in flint glass = 3*10^8 / 1.66 = 1.807*10^8 m/s

b) speed of light in water = 3*10^8 / 1.33 = 2.255 *10^8 m/s

c) speed of light in cubic zirconia = 3*10^8 / 2.2 = 1.363*10^8 m/s